OpenStudy (unklerhaukus):

Are there some more general sets i can add to this: \[\mathbb{N}\subset\mathbb{Z}\subset\mathbb{Q}\subset\mathbb{R}\subset\mathbb{C}\]

OpenStudy (mani_jha):

What do Q and C stand for?

OpenStudy (unklerhaukus):

Q is the ration numbers C is the complex numbers

OpenStudy (unklerhaukus):

rational*

OpenStudy (anonymous):

i think it's enough

OpenStudy (anonymous):

But there is D

OpenStudy (unklerhaukus):

what is D?

OpenStudy (anonymous):

D is the decimal numbers.

OpenStudy (anonymous):

A decimal number is a number with a decimal expansion limited, that is to say, a number which is written in any quantity, but finite, numbers after the decimal point.

OpenStudy (unklerhaukus):

ah i did not know that before, i guess they are between the integers Z and the Rational numbers Q \[\mathbb{N}\subset\mathbb{Z}\subset\mathbb{D}\subset\mathbb{Q}\subset\mathbb{R}\subset\mathbb{C}\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

but we généraly we talk about Q

OpenStudy (unklerhaukus):

yeah the D is a special Q where the denominator is 10^N

OpenStudy (mani_jha):

Hey, you could include W(whole numbers) between N and Z. They include all natural numbers and are included in the set of integers.

OpenStudy (unklerhaukus):

what is the difference between the whole numbers and the integers?

OpenStudy (mani_jha):

Whole numbers mean 0,1,2,3... and integers are ...-3,-2,-1,0,1,2,3...

OpenStudy (unklerhaukus):

ok then what is the difference between the natural numbers N and the whole numbers W?

OpenStudy (mani_jha):

Natural numbers include 1,2,3..They don't include 0, but whole numbers do.

OpenStudy (unklerhaukus):

ah ok, well i would define natural numbers to include 0 but not everybody does your definition would have \[\mathbb{N}\subset\mathbb{W} \subset\mathbb{N}\subset\mathbb{Z}\subset\mathbb{D}\subset\mathbb{Q}\subset\mathbb{R}\subset\mathbb{C}\]

OpenStudy (unklerhaukus):

\[\mathbb{N}\subset\mathbb{W} \subset\mathbb{N}\subset\mathbb{Z}\subset\mathbb{D}\subset\mathbb{Q}\subset\mathbb{R}\subset\mathbb{C} \subset\mathbb{H}\subset\mathbb{O}\subset\mathbb{S}\]